On August 1, Gilgen Brothers bought goods with a list price of $19,200, terms 2/10, n/30. The firm records purchases at invoice price, using the periodic inventory system. On August 5, Gilgen Brothers returned goods with a list price of $2,400 for credit.

If Gilgen Brothers paid the supplier the amount due on August 9, the appropriate entry would be:

Debit to Accounts Payable 18,464 Credit to Cash 18,464
Debit to Accounts Payable 19,200 Credit to Purchase Discounts 384 Cash 18,816
Debit to Accounts Payable 16,800 Credit to Cash 16,800
Debit to Accounts Payable 16,800 Purchases Discounts 336 Credit to Cash 16,464
Answer #1

List price of goods after return = $19,200 - 2400 = $16,800

Discount on balance = 16800 *0.02 = $336

The appropriate entry would be:

Debit to Accounts Payable 16,800 Purchases Discounts 336 Credit to Cash 16,464.